OpenStudy (anonymous):

A person jogs eight complete laps around a 400m track in a total time of 14.5 min.What is the average velocity?

OpenStudy (anonymous):

d = 400 m t = 14.5 min 14 min 30 sec total time will be = 870 sec it will take him to complete one track = 108.75 sec so velocity obtained per track will be v = d/t v = 400 /108.75 v = 3.7 m /s

OpenStudy (anonymous):

@leon549 They don't really learn a lot if you're just making th equestion instead of helping them step by step.

OpenStudy (anonymous):

average velocity is the displacement of a an object per unit time. displacment is the difference in the initial and final position in the given interval of time in which the average velocity has to be measured

OpenStudy (mrnood):

After 8 complete laps he is back where he started. Be sure you know the difference between average speed, and average velocity.

OpenStudy (mrnood):

@leon549 as @thomasker says: They learn little by being given the answer, and even less when it is not correct